OpenStudy (anonymous):

what are the steps of: 2m-13=8m+27

OpenStudy (anonymous):

m=14

OpenStudy (anonymous):

what are he steps

OpenStudy (anonymous):

oh, sorry...lol first, you have to get rid of the numbers next to "m", so subtract 2 from both "m's" and that will give you m-13=6m+27. Next you have to get rid of the -13, so add 13 to -13 (that will equal 0), and subtract 13 from 27, which will give you m=6m+14. Again, the "m" must be by itself so subtract that 6m from itself, and you have the answer of m=14.

OpenStudy (anonymous):

2m-8m=27+13 -6m=40 m=-40/6 = -20/3